OpenStudy (anonymous):

-How can one fifthx − 2 = one thirdx + 8 be set up as a system of equations? 5y − 5x = −10 3y − 3x = 24 5y − 5x = −10 3y + 3x = 24 5y + x = −10 3y + x = 24 5y − x = −10 3y − x = 24

OpenStudy (anonymous):

some one save the day this is the last question for me :/

OpenStudy (aravindg):

Last question for?

OpenStudy (anonymous):

homework

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

Can you use equation editor? one fifthx − 2 = one thirdx + 8

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }x-2=\frac{ 1 }{ 3 }x+8\]

OpenStudy (anonymous):

@AravindG i used it

OpenStudy (jessica_leon93):

@iPwnBunnies help her